OpenStudy (anonymous):

6(x+4)=2(x+5)????? :(

OpenStudy (anonymous):

???????:(

OpenStudy (blurbendy):

Distribute the 6 and 2, then get the x terms on the same side and solve for x

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

distributive property.

OpenStudy (anonymous):

6x + 24 = 2x + 10

OpenStudy (anonymous):

x= -7/2?

OpenStudy (anonymous):

\[6(x+4)=2(x+5)\]Distributive property:\[6x+24=2x+10\]Subtract \(2x\) from both sides (Subtraction Property of Equality):\[6x\color{blue}{-2x}+24=2x\color{blue}{-2x}+10\]\[4x+24=10\]Subtract \(10\) from both sides (Subtraction Property of Equality):\[4x+24\color{blue}{-24}=10\color{blue}{-10}\]\[4x=-14\]Divide both sides by \(4\) (Division Property of Equality):\[{4x\over\color{blue}{4}}={-14\over\color{blue}{4}}\]\[x=-{14\over4}\]Simplify:\[\large{x=-{7\over2}}\]